# Resolving UnderEdit & OverEdit with Iterative & Neighbor-Assisted Model Editingthanks: Preprint under-review

### v2

@@ 5.3 Analysis: How effective is the stopping criterion?
We tested two alternate stopping criteria to the proposed stopping criteria $|\Delta p_{k}|\leq 1$. The first is that $|\Delta p_{k}|$ should monotonically decrease i.e., $|\Delta p_{k+1}|<|\Delta p_{k}|$, otherwise stop and use $\theta_{k}$. The second is to stop when the difference in perplexity between consecutive iterations, i.e after Spread stage, is small, i.e., $\Delta_{p2}=|p(\theta_{k+1},h^{l_{c}}_{t,k+1})-p(\theta_{k},h^{l_{c}}_{t,k})|\leq 1$. We found our proposed criteria to be most the most effective in these experiments as shown in Table 4. Moreover, the second criteria suffered with two major drawbacks—a) it always needed at least two iterations to terminate and b) it always took one extra iteration longer than the proposed stopping criteria.
@@ 8 Limitations
The results shows that our proposed iterative and neighbor-assisted model editing approaches are highly effective resolving UnderEdit and OverEdit, respectively. However, we did notice some trade-offs where the former negatively impacted specificity and later generalization. We believe, these trade-offs stem from the fundamental challenge faced by direct model editing methods where LLM parameters are shared across different types of stored knowledge and currently no method exits to isolate parameters related to a knowledge. Our experiments and results highlight these challenges and encourages the research community to explore further resolving these challenges. So, we would recommend the adopters of our methods to prioritize between specificity or generalization depending on the application-specific requirements.
@@ Appendix A Locate-and-Edit Algorithms
As discussed in Section 2, model editing algorithms operate on the hypothesis that updating the final MLP parameters is sufficient to increase the likelihood of a new object $o^{*}$ over the original object $o$ when presented with a (subject, relation) pair $x=(s,r)$ as input to the LLM. Specifically, the final MLP weight matrix $W$ functions as a linear associative memory that stores a key-value mapping $[k,v]$ [^4] Distinct from key-value pairs in attention mechanismsMeng et al. (2023a). Here, the key encodes the last subject token, while the value represents the relation-object pair $(r,o)$ as a property of the subject $s$. Within the transformer architecture, the key corresponds to the output of the first fully connected MLP layer, whereas the value corresponds to the output of the second fully connected MLP layer $z$, as shown in Figure2. This key-value mapping $[k,v]$ is derived by computing the inner product between the key $k$ and the final MLP weight matrix $W$ as $Wk\approx v$.
Model editing involves modifying a batch of M desired edits, represented as D=$=\{(s_{m},r_{m},o_{m},o^{*}_{m})\}_{m=1}^{M}$, which translates to inserting $M$ new key-value pairs $[K_{M},V_{M}]$ by updating the final MLP weights $W$ at the causal layers $\{l_{1},\ldots,l_{c}\}$.
@@
Meng et al. (2023b) operates under the assumption that factual edits can be made by modifying the final MLP weight matrix $W^{l}$ at each causal layer $l$ with a small update $\Delta^{l}$, yielding new weights $W^{l^{*}}=W^{l}+\Delta^{l}$. The goal is to insert $M\gg 1$ new key-value mappings $[K_{M},V_{M}]$ while preserving $E$ existing mappings $[K_{E},V_{E}]$, where $K_{E}=[k_{e}]_{e=1}^{E}$ and $V_{E}=[v_{e}]_{e=1}^{E}$ denote the pre-existing keys and values.
@@
In practice, the residual $R^{l}$ is computed using the hidden state $\bar{h}^{l_{c}}_{t}$ obtained in the Optimization stage. For full derivation and further implementation details, see Meng et al. (2023b).
@@
Meng et al. (2023a) is the predecessor of MEMIT and adopts a more constrained approach to model editing. Unlike MEMIT, which apply updates across multiple causal layers, ROME assumes that new knowledge can be fully integrated into a single transformer layer $l_{*}$. It follows the same two-stage structure but focuses exclusively on editing one final MLP weight matrix $W^{l_{*}}$.
@@
The resulting solution allows for a rank-preserving edit that satisfies the new constraint while minimizing distortion to previously stored mappings. For a complete derivation, see Meng et al. (2023a).
@@ B.2 Neighbor-assisted model editing
This approach also introduced an additional constraint on the edit data points: each data point must have at least two neighboring knowledge samples—one for editing and the others for evaluation. Unfortunately, the ZsRE dataset contains only one neighboring knowledge sample per data point, restricting us to the COUNTERFACT dataset. Even within COUNTERFACT, only a limited number of samples met the required constraints. The number of qualifying samples varied depending on the model, as shown in Table 3. In neighbor-assisted model editing, exactly one neighboring example was used for each edit. Additionally, an average of 2, 4, and 4 neighboring examples (as reported in the Table 5) were used for evaluating each edit. Generating multiple neighbor knowledge using ChatGPT (Meng et al., 2023a) is straightforward and convenient. However, scaling this to all possible neighbors for a single piece of knowledge is a highly challenging research problem that falls beyond the scope of our paper.
Model Average evaluation neighboring samples (ceil)
GPT-2 XL 2
GPT-J 4
Llama-2-7B 4
@@
Llama 2 (7B) ROME COUNTERFACT A100 (80GB)

### v3

@@ 5.3 Analysis: How effective is the stopping criterion?
We tested two alternate stopping criteria to the proposed stopping criteria $|\Delta p_{k}|\leq 1$. The first is that $|\Delta p_{k}|$ should monotonically decrease i.e., $|\Delta p_{k+1}|<|\Delta p_{k}|$, otherwise stop and use $\theta_{k}$. The second is to stop when the difference in perplexity between consecutive iterations, i.e after Spread stage, is small, i.e., $\Delta_{p2}=|p(\theta_{k+1},h^{l_{c}}_{t,k+1})-p(\theta_{k},h^{l_{c}}_{t,k})|\leq 1$. We found our proposed criteria to be most the most effective in these experiments as shown in Table 4.
@@ 8 Limitations
We acknowledge the persistent trade-off between efficacy/generalization and specificity in locate-and-edit algorithms, which our methods do not eliminate. We believe, these trade-offs reflect a fundamental challenge of direct model editing: LLM parameters are shared across diverse knowledge types, and no existing method to the best of our knowledge can fully isolate the parameters tied to a specific fact. Our findings highlight this limitation and underscore the need for future research into more precise editing.
Our evaluation is limited to structured factual triplets, a constraint inherited from current locate-and-edit algorithms. Editing unstructured or free-form text remains challenging because locating target knowledge is non-trivial. We acknowledge this broader methodological limitation and view support for unstructured inputs as an important direction for future work.
@@
9 Ethical Statement
Our methods strengthen locate-and-edit model editing for factual updates. The same techniques could be used to insert false or misleading facts; we do not endorse such use. Similar to the general use of LLMs, models edited with our approach can still produce incorrect information or hallucinated content, so such system should be used with caution.
Acknowledgments
This research was supported in part by the University of Pittsburgh Center for Research Computing and Data, RRID:SCR_022735, through the resources provided. Specifically, this work used the H2P cluster, which is supported by NSF award number OAC-2117681. Bhiman was partly funded by a grant from Pitt Cyber Program [^4] https://www.cyber.pitt.edu/about. We are also grateful to Michael Boratko, the Pitt NLP group, and the anonymous reviewers for their constructive feedback and suggestions.
@@ Appendix A Locate-and-Edit Algorithms
As discussed in Section 2, model editing algorithms operate on the hypothesis that updating the final MLP parameters is sufficient to increase the likelihood of a new object $o^{*}$ over the original object $o$ when presented with a (subject, relation) pair $x=(s,r)$ as input to the LLM. Specifically, the final MLP weight matrix $W$ functions as a linear associative memory that stores a key-value mapping [k,v][k,v] [^5] Distinct from key-value pairs in attention mechanismsMeng et al. (2022). Here, the key encodes the last subject token, while the value represents the relation-object pair $(r,o)$ as a property of the subject $s$. Within the transformer architecture, the key corresponds to the output of the first fully connected MLP layer, whereas the value corresponds to the output of the second fully connected MLP layer $z$, as shown in Figure2. This key-value mapping $[k,v]$ is derived by computing the inner product between the key $k$ and the final MLP weight matrix $W$ as $Wk\approx v$.
[…]
@@
Meng et al. (2023) operates under the assumption that factual edits can be made by modifying the final MLP weight matrix $W^{l}$ at each causal layer $l$ with a small update $\Delta^{l}$, yielding new weights $W^{l^{*}}=W^{l}+\Delta^{l}$. The goal is to insert $M\gg 1$ new key-value mappings $K_{M},V_{M}]$ while preserving $E$ existing mappings $K_{E},V_{E}]$, where $K_{E}=[k_{e}]_{e=1}^{E}$ and $V_{E}=[v_{e}]_{e=1}^{E}$ denote the pre-existing keys and values.
@@
In practice, the residual $R^{l}$ is computed using the hidden state $\bar{h}^{l_{c}}_{t}$ obtained in the Optimization stage. For full derivation and further implementation details, see Meng et al. (2023).
@@
Meng et al. (2022) is the predecessor of MEMIT and adopts a more constrained approach to model editing. Unlike MEMIT, which apply updates across multiple causal layers, ROME assumes that new knowledge can be fully integrated into a single transformer layer $l_{*}$. It follows the same two-stage structure but focuses exclusively on editing one final MLP weight matrix $W^{l_{*}}$.
@@
The resulting solution allows for a rank-preserving edit that satisfies the new constraint while minimizing distortion to previously stored mappings. For a complete derivation, see Meng et al. (2022).
@@ B.1 Iterative model editing
"prompt": "The original language of Face Dances was",
"subject": "Face Dances",
"ground_truth": "English",
"target_new": "Italian",
"locality": {
"neighborhood": {
"prompt": [
"The original language of Ghost Rider is",
"The original language of 42nd Street was",
"The original language of The Fox and the Hound was",
"The language of Titanic is"
"ground_truth": [
"English",
"English",
"English",
"English"
"neighborhood_samples": [
"subject": "Star Wars: Episode III 2013 Revenge of the Sith",
"prompt": "The original language of {} was"
Figure 4: COUNTERFACT data sample for neighbor-assited model editing.
Model # Edited Examples Neighborhood Examples / Edit Average evaluation neighboring samples (ceil) / Edit
GPT-2 XL 739 1 2
GPT-J 690 1 4
Llama-2-7B 1340 1 4
@@ B.2 Neighbor-assisted model editing
This approach also introduced an additional constraint on the edit data points: each data point must have at least two neighboring knowledge samples—one for editing and the others for evaluation. Unfortunately, the ZsRE dataset contains only one neighboring knowledge sample per data point, restricting us to the COUNTERFACT dataset. Even within COUNTERFACT, only a limited number of samples met the required constraints. The number of qualifying samples varied depending on the model, as shown in Table 5. For clarity, we have added an illustrative example drawn from the COUNTERFACT dataset (Figure 4). In our neighbor-assisted model-editing setup, each edit consists of (i) a prompt—formed by the subject + relation pair—and (ii) a target_new (the new object). In addition, a single neighborhood_sample is always provided to guide the edit. Edit specificity is then assessed with the metric locality[neighborhood][prompt], whose neighborhood size varies as reported in Table 5. Therefore, in our neighbor-assisted model editing, exactly one neighboring example was used to assist the edit. Additionally, an average of 2, 4, and 4 neighboring examples (as reported in the Table 5) were used for evaluating each edit. Generating multiple neighbor knowledge using ChatGPT (Meng et al., 2022) is straightforward and convenient. However, scaling this to all possible neighbors for a single piece of knowledge is a highly challenging research problem that falls beyond the scope of our paper.
@@
Llama 2
ROME COUNTERFACT A100 (80GB)